# The algebraic and geometric classification of

### v1

@@ 1.3.1. Compatible pre-Lie algebras defined on $\mathcal{C}_{01}$
  • If $\alpha_{3}=0$, we obtain the parametric family:
    $\mathcal{C}_{10}^{\alpha,\beta}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.\qquad$
@@
1.3.2. Compatible pre-Lie algebras defined on $\mathcal{C}_{02}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{02},\mathcal{C}_{02})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{02}$ are of the form:
{$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2}.$\end{tabular}\right.\qquad$
Then we have the following cases:
  • If $\alpha_{1}\neq\alpha_{3}$, then choose $\nu=\frac{\alpha_{2}}{\alpha_{1}-\alpha_{3}}$ and obtain the parametric family:
    $\mathcal{C}_{11}^{\alpha,\beta\neq\alpha}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.\qquad$
@@ 1.3.2. Compatible pre-Lie algebras defined on $\mathcal{C}_{02}$
  • If $\alpha_{1}=\alpha_{3}$, we obtain the parametric family:
    $\mathcal{C}_{12}^{\alpha,\beta}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.\qquad$
@@
1.3.3. Compatible pre-Lie algebras defined on $\mathcal{C}_{03}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{03},\mathcal{C}_{03})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{03}$ are of the forms:
@@ 1.3.3. Compatible pre-Lie algebras defined on $\mathcal{C}_{03}$
    {$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2},$&$e_{2}\ast e_{1}=\alpha_{1}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1}+\alpha_{5}e_{2},$&$e_{1}\ast e_{2}=\alpha_{6}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{2},$&&&\\ $e_{1}\ast e_{1}=\alpha_{7}e_{1}+\alpha_{8}e_{2},$&$e_{1}\ast e_{2}=\alpha_{9}e_{1},$&$e_{2}\ast e_{1}=\alpha_{9}e_{1},$&$e_{2}\ast e_{2}=\alpha_{9}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{2},$&&&\\ $e_{1}\ast e_{1}=\alpha_{10}e_{1}+\alpha_{11}e_{2},$&$e_{1}\ast e_{2}=2\alpha_{10}e_{2},$&$e_{2}\ast e_{1}=\alpha_{12}e_{1}+\alpha_{10}e_{2},$&$e_{2}\ast e_{2}=2\alpha_{12}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{4}\alpha_{9}\alpha_{12}\neq 0$.
First, we consider the first form. Then we have the following cases:
  • If $\alpha_{3}\neq 0$, then choose $\xi=\alpha^{-1}_{3},\nu=-{\alpha_{2}}{\alpha_{3}^{-2}}$ and obtain the parametric family:
    $\mathcal{C}_{13}^{\alpha}:=\left\{\begin{tabular}[]{lll}$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.\qquad$
@@
  • If $\alpha_{3}=0,\alpha_{1}\neq 0$, then choose $\xi={\alpha^{-1}_{1}}$ and obtain the parametric family:
    $\mathcal{C}_{14}^{\alpha}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{1}=e_{2}.$\end{tabular}\right.\qquad$
@@
  • If $\alpha_{1}=\alpha_{3}=0$, then we obtain the parametric family:
    $\mathcal{C}_{15}^{\alpha}:=\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.\qquad$
@@
  • If $\alpha_{4}\neq\alpha_{6}$, then choose $\xi={\alpha^{-1}_{4}},\nu=\frac{\alpha_{5}}{\alpha_{4}-\alpha_{6}}$ and obtain the parametric family:
    $\mathcal{C}_{16}^{\alpha}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.$
@@
  • If $\alpha_{4}=\alpha_{6}$, then choose $\xi={\alpha^{-1}_{4}}$ and obtain the parametric family:
    $\mathcal{C}_{17}^{\alpha}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{1}\ast e_{2}=e_{2}.$\end{tabular}\right.$
@@
Third, we consider the third form. Then, we choose $\xi={\alpha^{-\frac{1}{2}}_{9}},\nu=-\frac{\alpha_{7}\alpha_{9}^{-\frac{3}{2}}}{2}$ and obtain the parametric family:
$\mathcal{C}_{18}^{\alpha}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{2},$&&&\\ $e_{1}\ast e_{1}=\alpha e_{2},$&$e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
@@
Finally, we consider the fourth form. Then, we choose $\xi=\alpha^{-\frac{1}{2}}_{12},$ $\nu=-{\alpha_{10}}{\alpha_{12}^{-\frac{3}{2}}}$ and obtain the parametric family:
$\mathcal{C}_{19}^{\alpha}:=\left\{\begin{tabular}[]{lll}$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=2e_{2}.$\end{tabular}\right.$
@@
1.3.4. Compatible pre-Lie algebras defined on $\mathcal{C}_{04}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{04},\mathcal{C}_{04})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{04}$ are of the form:
@@ 1.3.4. Compatible pre-Lie algebras defined on $\mathcal{C}_{04}$
    {$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha_{4}e_{1},$&$e_{2}\ast e_{1}=\alpha_{5}e_{1},$&$e_{2}\ast e_{2}=\alpha_{6}e_{1}+\alpha_{4}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{4}\neq 0$.
First, we study the first form. Then we have the following cases:
  • If $\alpha_{2}\neq 0$, then choose $\xi=\alpha_{2}$ and obtain the parametric family:
[…]
    =e2e1e1,
    =e2e1αe1, =e2e2+e1βe2.
[…]
@@
  • If $\alpha_{2}=0$, we obtain the parametric family:
    $\mathcal{C}_{21}^{\alpha,\beta}:=\left\{\begin{tabular}[]{ll}$e_{2}\cdot e_{1}=e_{1},$&\\ $e_{2}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.$
@@
  • If $\alpha_{6}\neq 0$, then choose $\xi=\alpha_{6}$ and obtain the parametric family:
    $\mathcal{C}_{22}^{\alpha\neq 0,\beta}:=\left\{\begin{tabular}[]{lll}$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha e_{1},$&$e_{2}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{2}=e_{1}+\alpha e_{2}.$\end{tabular}\right.$
@@
  • If $\alpha_{6}=0$, we obtain the parametric family:
    $\mathcal{C}_{23}^{\alpha\neq 0,\beta}:=\left\{\begin{tabular}[]{lll}$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha e_{1},$&$e_{2}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.$
@@
1.3.5. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{\alpha\neq 0,\frac{1}{2},1}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{05}^{\alpha},\mathcal{C}_{05}^{\alpha})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\alpha}$ are of the form:
{$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2}.$\end{tabular}\right.$
Then we have the following cases:
  • If $\alpha_{2}\neq 0$, then choose $\xi=\alpha_{2}$ and obtain the parametric family:
[…]
    =e1e1e1, =e1e2αe2,
    =e1e1+βe1e2, =e1e2γe2.
[…]

### v2

@@ 2.3.1 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{01}}$
  • If $\alpha_{3}=0$, we obtain the parametric family
    $\mathcal{C}_{10}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2}.$
@@
2.3.2 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{02}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{02},\mathcal{C}_{02})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{02}$ are of the form $e_{1}\cdot e_{1}=e_{1}+e_{2}$, $e_{1}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$. Then we have the following cases:
  • If $\alpha_{1}\neq\alpha_{3}$, then choose $\nu=\frac{\alpha_{2}}{\alpha_{1}-\alpha_{3}}$ and obtain the parametric family
    $\mathcal{C}_{11}^{\alpha,\beta\neq\alpha}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\beta e_{2}.$
@@ 2.3.2 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{02}}$
  • If $\alpha_{1}=\alpha_{3}$, we obtain the parametric family
    $\mathcal{C}_{12}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{1}\ast e_{2}=\alpha e_{2}.$
@@
2.3.3 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{03}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{03},\mathcal{C}_{03})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{03}$ are of the following forms:
@@ 2.3.3 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{03}}$
    $e_{1}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}+\alpha_{5}e_{2}$, $e_{1}\ast e_{2}=\alpha_{6}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{7}e_{1}+\alpha_{8}e_{2}$, $e_{1}\ast e_{2}=\alpha_{9}e_{1}$, $e_{2}\ast e_{1}=\alpha_{9}e_{1}$, $e_{2}\ast e_{2}=\alpha_{9}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{10}e_{1}+\alpha_{11}e_{2}$, $e_{1}\ast e_{2}=2\alpha_{10}e_{2}$, $e_{2}\ast e_{1}=\alpha_{12}e_{1}+\alpha_{10}e_{2}$, $e_{2}\ast e_{2}=2\alpha_{12}e_{2}$.
We may assume $\alpha_{4}\alpha_{9}\alpha_{12}\neq 0$. First, we consider the first form. Then we have the following cases:
  • If $\alpha_{3}\neq 0$, then choose $\xi=\alpha^{-1}_{3}$, $\nu=-{\alpha_{2}}{\alpha_{3}^{-2}}$ and obtain the parametric family
    $\mathcal{C}_{13}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2}.$
@@
  • If $\alpha_{3}=0,\alpha_{1}\neq 0$, then choose $\xi={\alpha^{-1}_{1}}$ and obtain the parametric family
    $\mathcal{C}_{14}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1}+\alpha e_{2},\quad e_{2}\ast e_{1}=e_{2}.$
@@
  • If $\alpha_{1}=\alpha_{3}=0$, then we obtain the parametric family
    $\mathcal{C}_{15}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2}.$
@@
  • If $\alpha_{4}\neq\alpha_{6}$, then choose $\xi={\alpha^{-1}_{4}}$, $\nu=\frac{\alpha_{5}}{\alpha_{4}-\alpha_{6}}$ and obtain the parametric family
    $\mathcal{C}_{16}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2}.$
@@
  • If $\alpha_{4}=\alpha_{6}$, then choose $\xi={\alpha^{-1}_{4}}$ and obtain the parametric family
    $\mathcal{C}_{17}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1}+\alpha e_{2},\quad e_{1}\ast e_{2}=e_{2}.$
@@
Third, we consider the third form. Then, we choose $\xi={\alpha^{-\frac{1}{2}}_{9}}$, $\nu=-\frac{\alpha_{7}\alpha_{9}^{-\frac{3}{2}}}{2}$ and obtain the parametric family
$\mathcal{C}_{18}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
@@
Finally, we consider the fourth form. Then, we choose $\xi=\alpha^{-\frac{1}{2}}_{12}$, $\nu=-{\alpha_{10}}{\alpha_{12}^{-\frac{3}{2}}}$ and obtain the parametric family
$\mathcal{C}_{19}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=2e_{2}.$
@@
2.3.4 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{04}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{04},\mathcal{C}_{04})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{04}$ are of the following forms:
@@ 2.3.4 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{04}}$
    $e_{2}\cdot e_{1}=e_{1}$, $e_{1}\ast e_{2}=\alpha_{4}e_{1}$, $e_{2}\ast e_{1}=\alpha_{5}e_{1}$, $e_{2}\ast e_{2}=\alpha_{6}e_{1}+\alpha_{4}e_{2}$.
We may assume $\alpha_{4}\neq 0$. First, we study the first form. Then we have the following cases:
  • If $\alpha_{2}\neq 0$, then choose $\xi=\alpha_{2}$ and obtain the parametric family
    $\mathcal{C}_{20}^{\alpha,\beta}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{2}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=e_{1}+\beta e_{2}.$
@@
  • If $\alpha_{2}=0$, we obtain the parametric family
    $\mathcal{C}_{21}^{\alpha,\beta}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{2}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=\beta e_{2}.$
@@
  • If $\alpha_{6}\neq 0$, then choose $\xi=\alpha_{6}$ and obtain the parametric family
    $\mathcal{C}_{22}^{\alpha\neq 0,\beta}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{1},\quad e_{2}\ast e_{1}=\beta e_{1},\quad e_{2}\ast e_{2}=e_{1}+\alpha e_{2}.$
@@
  • If $\alpha_{6}=0$, we obtain the parametric family
    $\mathcal{C}_{23}^{\alpha\neq 0,\beta}\colon\ e_{2}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{1},\quad e_{2}\ast e_{1}=\beta e_{1},\quad e_{2}\ast e_{2}=\alpha e_{2}.$
@@
2.3.5 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{\alpha\not=0,\frac{1}{2},1}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{05}^{\alpha},\mathcal{C}_{05}^{\alpha})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\alpha}$ are of the form $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=\alpha e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$. Then we have the following cases:
  • If $\alpha_{2}\neq 0$, then choose $\xi=\alpha_{2}$ and obtain the parametric family
    $\mathcal{C}_{24}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{1}\ast e_{1}=\beta e_{1}+e_{2},\quad e_{1}\ast e_{2}=\gamma e_{2}.$